### v3

# RsGCN: Rescaling Enhances Generalization of GCNs for Solving Scalable Traveling Salesman Problems

Abstract
Neural traveling salesman problem (TSP) solvers face two critical challenges: poor generalization for scalable TSPs and high training costs. To address these challenges, we propose a new Rescaling Graph Convolutional Network (RsGCN). Focusing on the scale-dependent features (i.e., features varied with problem scales) related to nodes and edges that influence the sensitivity of GCNs to the problem scales, a Rescaling Mechanism in RsGCN enhances the generalization capability by (1) rescaling adjacent nodes to construct a subgraph with a uniform number of adjacent nodes for each node across various scales of TSPs, which stabilizes the graph message aggregation; (2) rescaling subgraph edges to adjust the lengths of subgraph edges to the same magnitude, which maintains numerical consistency. In addition, an efficient training strategy with a mixed-scale dataset and bidirectional loss is used in RsGCN. To fully exploit the heatmaps generated by RsGCN, we design an efficient post-search algorithm termed Re2Opt, in which a reconstruction process based on adaptive weight is incorporated to help avoid local optima. Based on a combined architecture of RsGCN and Re2Opt, our solver achieves remarkable generalization and low training cost: with only 3 epochs of training on the mixed-scale dataset containing instances with up to 100 nodes, it can be generalized successfully to 10K-node instances without any fine-tuning. Extensive experiments demonstrate our state-of-the-art performance across uniform distribution instances of 9 different scales from 20 to 10K nodes and 78 real-world instances from TSPLIB, while requiring the fewest learnable parameters and training epochs among neural competitors.
1 Introduction
Background
The traveling salesman problem (TSP), as a typical combinatorial optimization problem, has been extensively studied in the literature. Although classical solvers like Concorde [1] and LKH [7] achieve strong performance across various problem scales, their designs highly rely on expert-crafted heuristics. The development of deep learning has spawned numerous neural TSP solvers with various learning paradigms (e.g., Supervised Learning (SL), Reinforcement Learning (RL)) and architectural foundations (e.g., Graph Convolutional Networks (GCNs) [6, 9, 12, 18, 19, 27], Transformers [5, 15, 16, 17, 22, 24, 33], and Diffusion Models [18, 19, 27]). Herein, a typical framework is using neural networks to generate heatmaps and then incorporate post-search algorithms to obtain solutions with the guidance of heatmaps [6, 18, 19, 27]. Monte Carlo Tree Search [6], owing to its promising search efficiency, is frequently employed in the post-search guided by heatmaps.
Challenges
Although prior work has shown neural networks’ substantial potential in solving TSPs, it also reveals critical limitations in solution generalization [11]. Specifically, pre-trained models cannot perform well for new problem scales, and thus further fine-tuning with new-scale instances is necessary to transfer pre-trained models to the new problem scales. Existing approaches to enhance generalization include: (1) enhancing solution diversity by generating multiple heatmaps [18, 19, 27] or designing specific post-search [5, 6, 8, 9], and (2) decomposing large-scale problems into small-scale sub-problems for partitioned solving [6, 17, 22, 24, 33]. Although these approaches achieve modest improvements, they suffer from prohibitive computational overhead. Regarding (1), generating additional heatmaps is computationally intensive for GPUs, significantly diminishing the practicality. Furthermore, existing post-search algorithms exhibit insufficient capability in narrowing down the search space and thus result in limited search efficiency. As for (2), the decomposition of large-scale problems leads to a partial loss of global perspective and thus weakens the overall optimization effectiveness, while the design of sub-problems partitioning and recombination introduces significant complexity.
Our Contributions
Based on our research analysis, since neural networks are sensitive to numerical values, the scale-dependent features related to nodes and edges in scalable TSPs are essential factors impairing generalization. In this paper, we propose a new GCN termed RsGCN, which incorporates a Rescaling Mechanism to adaptively rescale the scale-dependent features. As shown in Figure 1, the Rescaling Mechanism (1) rescales adjacent nodes to stabilize the graph message aggregation, which constructs a subgraph for each node by k-Nearest Neighbor Selection to ensure each node has a uniform number of adjacent nodes regardless of instance scale; (2) rescales subgraph edges to maintain numerical consistency, which adjusts the length of edges in subgraphs to the same magnitude by Uniform Unit Square Projection. With the Rescaling Mechanism, RsGCN can learn universal patterns by training on only small-scale instances. Thus, RsGCN can be effortlessly generalized to large-scale TSPs and generate high-quality heatmaps to guide post-search for scalable TSPs. In addition, rethinking the limitations of existing post-search algorithms, we design a new post-search algorithm termed Reconstruction-2Opt (Re2Opt), which utilizes 2Opt as the basic optimizer and performs reconstruction based on adaptive weights to robustly escape from local optima.
According to extensive experimental studies, the combined architecture of RsGCN and Re2Opt shown in Figure 1 achieves state-of-the-art performance on both uniform distribution and real-world instances. Our contributions are summarized as follows:
  1. 1.
    Rescaling Mechanism for Enhancing Generalization: We propose a Rescaling Mechanism to rescale the number of adjacent nodes and the lengths of subgraph edges, enhancing GCN into RsGCN to significantly improve generalization capability in solving scalable TSPs. With the Rescaling Mechanism, RsGCN requires significantly fewer learnable parameters and training epochs compared with other neural solvers and only uses a mixed-scale dataset containing instances with up to 100 nodes for training, while the obtained model can be effectively generalized to large-scale instances (e.g., 10K-node instances) without any fine-tuning.
  2. 2.
    Efficient Post-Search for Leveraging Heatmaps: To sufficiently utilize the informative heatmaps generated by RsGCN, we design a new Re2Opt algorithm for post-search, which incorporates reconstruction to help escape from local optima.
  3. 3.
    Ordered Heatmaps for Visual Analysis: We introduce ordered heatmaps to facilitate the visual analysis of the heatmaps’ quality and help validate the effectiveness of the Rescaling Mechanism.
Figure 1: The architecture of RsGCN with heatmap-guided post-search Re2Opt.
2 Related Work
Based on the differences in the way of constructing solutions, existing neural TSP solvers can be divided into two categories: rule-based constructive solvers and learning-based constructive solvers.
2.1 Rule-Based Constructive (RC) Solvers
In RC solvers [12, 6, 9, 27, 18, 19], the trained neural networks output an informative heatmap that indicates promising connections between nodes. Subsequently, the heatmap is used to guide the post-search algorithm to construct the TSP tour. In other words, neural networks are only involved in the generation of the heatmap and do not directly obtain valid solutions for TSPs. The encoders of RC solvers are typically based on GCN, while MLP [12, 6, 9] or Diffusion [27, 18, 19] are employed as decoders.
In the post-search with heatmaps, simple algorithms (e.g., greedy search and beam search) are used to generate TSP tours as initial states [12, 9, 27, 19] and 2Opt [4] algorithm is used to further optimize the generated TSP tours [27, 19]. In addition, diffusion-based models generate multiple heatmaps with different random seeds and then conduct greedy search to obtain multiple solutions. To perform search with a broader range based on the heatmap, Monte Carlo Tree Search (MCTS) [6] is proposed based on a Markov decision process to perform k-Opt operations. In general, both high-quality heatmaps and efficient post-search algorithms are crucial in RC solvers.
2.2 Learning-Based Constructive (LC) Solvers
LC solvers [15, 16, 5, 24, 22, 33, 17] take the current solution state as the input of neural networks and output the complete TSP tours. Such a solving process usually involves divide-and-conquer and inference within multiple iterations, which are mainly learned by neural networks. For example, H-TSP [24] employs a hierarchical approach with lower-level and upper-level models, in which the lower-level model solves sub-problems while the upper-level model combines the solutions of sub-problems into a complete TSP tour. DRHG [17] breaks edges within a cluster and constructs a sub-problem based on a hyper-graph structure before optimization. Transformer-based models with autoregressive construction have been effectively applied in LC solvers. Note that although LC solvers also use some simple rule-based search algorithms to construct solutions, most of the computational overhead occurs in the autoregressive and iterative process of neural networks.
2.3 Rethinking Complexity and Generalization
The heatmaps obtained by RC solvers do not show sufficient generalization capability in scalable TSPs and thus need to be fine-tuned with multiple new-scale instances to adapt to scalable TSPs, which increases the training cost. To enhance solution diversity, the diffusion-based decoder is used [27, 18, 19], which generates multiple heatmaps with multiple random seeds. However, the multi-step denoising process in the diffusion-based decoder is computationally intensive during training and inference. In addition, although some post-search algorithms such as MCTS achieve promising performance, their hyperparameters should be tuned for different scales of TSPs, which relies on expert knowledge and weakens the generalization capability. A recent study [31] has also questioned the role of heatmaps in guiding MCTS because MCTS can perform better without heatmaps generated by neural networks.
LC solvers typically conduct divide-and-conquer to decompose a large-scale TSP into several small-scale sub-problems [24, 22, 33, 17], and thus they have a certain capability of generalization for scalable TSPs. However, the optimization of each sub-problem loses the global perspective and may easily obtain a local optimum. If no problem decomposition is conducted or the scale of sub-problems is enlarged, which can enhance the global search capability, the computational complexity of LC solvers will extensively increase and result in unacceptable solution time overhead. In addition, the decomposition of sub-problems also relies on expert knowledge. In [24, 33], a hierarchical problem-solving framework is employed, which combines multiple models trained on different scales of TSP datasets to enhance the generalization capability. However, such a framework is complex and requires expensive computational costs.
To enhance the generalization capability while reducing the complexity and computational cost, we incorporate a Rescaling Mechanism into GCN and propose a new RsGCN, which learns universal patterns across various problem scales by rescaling the scale-dependent features related to nodes and edges in TSPs. With only 3 training epochs on a mixed-scale dataset (including instances with 20, 30, 50, and 100 nodes), RsGCN can be successfully generalized to 10K-node instances without any fine-tuning. An experimental investigation is also conducted to validate the efficient guidance provided by the heatmaps of RsGCN. In addition, a new post-search algorithm Re2Opt is proposed, which does not require dedicated hyperparameter tuning for different scales of TSPs.
3 Preliminaries
Problem Definition
In TSPs, nodes are located in a 2-dimensional Euclidean space with undirected edges between nodes. For an $n$-node instance, $\mathcal{X}_{n}=\{x_{1},x_{2},\dots,x_{n}\}$ denotes its set of node coordinates, where $x_{i}=(a_{i},b_{i})$. A solution for TSPs, i.e., a TSP tour, is a closed loop that traverses all nodes exactly once and is typically encoded by a permutation $\Pi_{n}=\{\pi_{1},\pi_{2},\dots,\pi_{n}\}$. The length of the tour $\Pi_{n}$, denoted by $L(\Pi_{n})$, is calculated as follows, and the optimization objective of TSP is finding a permutation with the minimum tour length.
$L(\Pi_{n})=\sum_{i=1}^{n-1}(\|x_{\pi_{i}}-x_{\pi_{i+1}}\|^{2})+\|x_{\pi_{n}}-x_{\pi_{1}}\|^{2}$ (1)
Heatmap Representation
GCNs output a heatmap $\boldsymbol{H}\in\mathbb{R}^{n\times n}$ with input of an $n$-node instance, where $H_{i,j}$ represents the heat between $x_{j}$ and $x_{i}$, i.e., the probability that $x_{j}$ is adjacent to $x_{i}$ in the generated TSP tour. To facilitate observation and analysis, we define ordered heatmaps that are sorted according to the optimal tour. Details of ordered heatmaps are shown in Appendix A.
4 Methodology
4.1 Rescaling Mechanism for GCN (RsGCN)
GCNs have powerful feature representation capability to generate high-quality heatmaps for fixed-scale TSPs. However, previous studies [11, 27] have shown that GCNs exhibit limited generalization capability for scalable TSPs. We find that the primary issue stems from GCNs’ sensitivity to scale-dependent features, including (1) the number of adjacent nodes, and (2) the measurement of distances from any given node to its adjacent nodes. Specifically, the scale-dependent features vary in different scales of TSPs and thus weaken the generalization capability. To enhance the generalization capability for scalable TSPs, the proposed RsGCN incorporates the Rescaling Mechanism to rescale the scale-dependent features and help learn universal patterns across TSPs with different scales, which consists of rescaling adjacent nodes and rescaling subgraph edges as follows.
4.1.1 Rescaling Adjacent Nodes
The connections between nodes in TSP form fully connected graphs, implying that the adjacent nodes of each node increase as the total number of nodes increases. This not only leads to a significant increase in computational complexity but also poses great challenges to the robustness of GCNs’ representations when performing message aggregation on dense graphs of various scales.
To address this issue, we employ k-Nearest Neighbor Selection to rescale the number of adjacent nodes as $k_{1}$ for each node, thereby each node forms a subgraph including itself and its top $k_{1}$ nearest neighbors. $k_{1}$ is fixed and thus the nodes in TSPs with various scales have subgraphs with a uniform number of adjacent nodes, helping learn universal patterns from instances with different scales. In addition, it can also achieve graph sparsification [20] that enhances GCNs by retaining promising candidate nodes and stabilizing message aggregation.
Specifically, for an $n$-node instance with node set $\mathcal{X}_{n}=\{x_{1},x_{2},\dots,x_{n}\}$, the subgraph set is denoted by $\boldsymbol{\mathcal{N}}(\mathcal{X}_{n}|k_{1})=\{\mathcal{N}_{1}^{k_{1}},\mathcal{N}_{2}^{k_{1}},\dots,\mathcal{N}_{n}^{k_{1}}\}$, where $\mathcal{N}_{i}^{k_{1}}=\{x_{i}^{1},x_{i}^{2},\dots,x_{i}^{k_{1}}\}$ contains the $k_{1}$-nearest neighbors of $x_{i}$ (including $x_{i}$ itself) and represents $x_{i}$’s subgraph. $x_{i}^{j}=(a_{i}^{j},b_{i}^{j})\in\mathcal{X}_{n}$ denotes the coordinate of $x_{i}$’s $j$-th nearest neighbor. Note that after rescaling the adjacent nodes, the range of message aggregation is narrowed down and the complexity of message aggregation in GCNs is reduced from $O(n^{2})$ to $O(nk_{1})$.
4.1.2 Rescaling Subgraph Edges
In existing research, the coordinates of nodes in TSPs are typically normalized in a unit square space $[0,1]^{2}$ to uniform the range of node distribution and thus can facilitate the training of neural networks. However, such a normalization also influences the edge features in TSPs, which is neglected. In detail, with the same range of space, the distribution of nodes in an instance with more nodes is denser, which makes the distances between nodes smaller on average. Therefore, the edge features about distances between nodes vary in scalable TSPs, which limits the generalization capability of neural TSP solvers.
To deal with the above issue, the Uniform Unit Square Projection is proposed and incorporated into the Rescaling Mechanism. For a subgraph of a node $x_{i}$, the overall principle is to rescale the coordinates of nodes in $x_{i}$’s subgraph and obtain a transformed subgraph in which the coverage of $x_{i}$ and its neighbors is maximized in the unit square space.
In detail, we first acquire the maximum and minimum coordinate values of nodes on the two axes from $\mathcal{N}_{i}^{k_{1}}$, denoted by $a_{i}^{\max}$, $a_{i}^{\min}$, $b_{i}^{\max}$, and $b_{i}^{\min}$, respectively. Then, we compute the scaling factor $\mu_{i}$ as $\mu_{i}=\frac{1}{\max(A_{i},B_{i})}$, where $A_{i}=a_{i}^{\max}-a_{i}^{\min}$ and $B_{i}=b_{i}^{\max}-b_{i}^{\min}$. Based on the above calculations, we rescale each node coordinate $x_{i}^{j}=(a_{i}^{j},b_{i}^{j})$ in $\mathcal{N}_{i}^{k_{1}}$ to obtain the rescaled coordinate $r(x_{i}^{j})$ according to Equation 2 and finally project $\mathcal{N}_{i}^{k_{1}}$ into $R(\mathcal{N}_{i}^{k_{1}})$ according to Equation 3. Herein, $\mu_{i}$ is used to maximize the coverage of node distribution in the unit square space while ensuring that the rescaled coordinate does not exceed the boundary of the unit square space.
$r(x_{i}^{j})=r[(a_{i}^{j},b_{i}^{j})]=((a_{i}^{j}-a_{i}^{\min})\times\mu_{i},\ (b_{i}^{j}-b_{i}^{\min})\times\mu_{i}),\ \forall j\in\{1,2,\dots,k_{1}\}$ (2)
$R(\mathcal{N}_{i}^{k_{1}})=\{r(x_{i}^{1}),r(x_{i}^{2}),\dots,r(x_{i}^{k_{1}})\},\ \forall i\in\{1,2,\dots,n\}$ (3)
After rescaling subgraph edges, we calculate the Euclidean distances from $x_{i}$ to its $k_{1}$-nearest neighbors based on $R(\mathcal{N}_{i}^{k_{1}}),\forall i\in\{1,2,\dots,n\}$. Consequently, the edge lengths in each subgraph from TSP instances with different scales are rescaled to the same magnitude, reducing the influence of distance measurement on the generalization capability in dealing with scalable TSPs. For intuitive comprehension, an example of the Uniform Unit Square Projection is shown in Figure 2. The leftmost figure represents the original distribution of the nodes in a subgraph, and the red rectangle denotes the subgraph frame constructed based on the marginal nodes of the subgraph. After the two steps illustrated in Figure 2, the rescaled subgraph is obtained and is shown in the rightmost figure. Note that each arrow represents a step, with the specific operation displayed above and the related variables displayed below the arrow. In addition, the effect of the Rescaling Mechanism is visualized and discussed in more detail in Appendix H.
Figure 2: An example of the Uniform Unit Square Projection for a subgraph.
4.1.3 Architecture of GCN
Due to space limitations, this section has been placed in Appendix B.
4.2 Training Strategy
Mixed-Scale Dataset
Since RsGCN acquires powerful generalization capability with the Rescaling Mechanism, a mixed-scale dataset is used to train RsGCN. In detail, the mixed-scale dataset includes one million TSP instances, in which the number of nodes varies in 20, 30, 50, and 100, and the proportion of these four scales is 1:2:3:4. On the one hand, the mixed-scale dataset effectively helps GCNs learn universal patterns across scalable TSPs, avoiding the pitfalls of narrow patterns specific to a single scale. On the other hand, the maximum scale in the mixed-scale dataset is only 100 nodes, and RsGCN can be efficiently generalized to 10K-node instances without any fine-tuning, which can significantly reduce the training cost.
Bidirectional Loss
The single-label cross-entropy with Softmax is widely used in the loss function of neural TSP solvers. However, previous work does not consider that each node is connected to exactly two other nodes in TSP tours. Accordingly, we employ a double-label binary cross-entropy with Sigmoid as the loss function. In detail, the two adjacent nodes of $x_{i}$ in the optimal tour $\hat{\Pi}$ serve as the two labels. The loss function $\mathcal{L}$ for $n$-node instances is presented as follows:
$\mathcal{L}$ $=-\frac{\sum_{x_{i}\in\mathcal{X}_{n}}\sum_{x_{j}\in\mathcal{X}_{n}}\mathcal{Z}_{i,j}}{n}$ (4)
$\mathcal{Z}_{i,j}$ =$=\begin{cases}\log(H_{i,j}+\varepsilon),&\text{if }x_{j}\text{ isadjacent to }x_{i}\text{ in }\hat{\Pi}\\ \log(1-H_{i,j}+\varepsilon),&\text{others}\end{cases}$ (5)
where the division by $n$ in Equation 4 aims to balance the loss for scalable TSPs, helping prevent the gradient updates from being biased towards larger-scale instances. In Equation 5, $H_{i,j}$ is the probability between $x_{i}$ and $x_{j}$ in the heatmap normalized by the Sigmoid function. $\varepsilon$ is a very small value used to avoid taking the logarithm of zero.
4.3 Efficient Post-Search – Re2Opt
An efficient post-search algorithm is essential to fully utilize heatmaps. Inspired by the k-Opt operations in MCTS, we propose a more efficient and robust search algorithm termed Reconstruction-2Opt (Re2Opt). The overall procedure of Re2Opt is shown in Figure 1, which first conducts State Initialization to obtain an initial tour and then iteratively conducts Reconstruction and 2Opt and Weight Enhancement for further optimization until the time budget is exhausted.
4.3.1 State Initialization
The initial tour is constructed by greedy selection. First, a node $x_{i}$ is randomly selected from $\mathcal{X}_{n}$. Second, a candidate set of the next node for $x_{i}$ is constructed, denoted by $\mathcal{C}_{i}$. In detail, $\mathcal{C}_{i}$ consists of the top $k_{2}$ hottest nodes for $x_{i}$, i.e., the top $k_{2}$ nodes with the highest probability between $x_{i}$ according to the generated heatmap. Third, the hottest node in $\mathcal{C}_{i}$ that has not been traversed yet is selected as the next node. If all nodes in $\mathcal{C}_{i}$ have been traversed but the TSP tour is still incomplete, the nearest untraversed node to $x_{i}$ is selected as the next node. Then, the edge between $x_{i}$ and the next node is adopted, and $x_{i}$ is updated to the selected next node. The second and third steps are conducted iteratively until all nodes are traversed. Finally, the initial tour is obtained.
Figure 3: An example of reconstruction in Re2Opt.
4.3.2 Reconstruction
Reconstruction aims to enhance the search diversity and escape from local optima. First, a node is randomly selected as the split node. There are two edges directly connected to the split node in the tour, and the one with a smaller weight is selected and removed to form an acyclic path. Herein, the weights of edges are stored in $\boldsymbol{Q}$, where $\boldsymbol{Q}_{i,j}$ represents the importance of the edge $x_{i}$-$x_{j}$. $\boldsymbol{Q}$ is initialized as an all-zero symmetric matrix and is updated during the 2Opt and Weight Enhancement. Then, taking the split node as the start node, a reconstruction action is conducted on the current acyclic path. In detail, a target node is selected from non-adjacent nodes of the start node based on the probability distribution $t_{i,j}=\frac{\boldsymbol{Q}_{i,j}}{\sum_{x_{h}\in\mathcal{C}_{i}}\boldsymbol{Q}_{i,h}}$, where $t_{i,j}$ represents the probability of $x_{j}$ being selected as the target node with $x_{i}$ as the start node. Finally, the acyclic path is updated by removing a directly connected edge of the target node and adding a new edge between the start node and the target node. Note that after a reconstruction action, the start node is also updated, and multiple reconstruction actions are conducted in a reconstruction process.
Take Figure 3 as an example to illustrate details. Initially, node 2 is the randomly selected split node. The two edges directly connected to node 2 are “2-1” and “2-3”. If the weight of “2-3” is not larger than that of “2-1”, i.e., $\boldsymbol{Q}_{2,1}\geq\boldsymbol{Q}_{2,3}$, “2-3” is removed from the tour and an acyclic path 2-1-5-4-3 is obtained. A reconstruction action is illustrated on the right side of Figure 3. Assuming node 4 is selected as the target node of the current start node (node 2), the edge between the target node and its previous adjacent node, i.e., the edge “5-4”, is removed. A new edge between the start node and the target node, i.e., the edge “2-4”, is added. After that, we can obtain a new acyclic path 5-1-2-4-3 with node 5 as the start node for the next reconstruction action.
The reconstruction process iteratively conducts reconstruction actions and will terminate until any one of the following three criteria is met: (1) The tour formed by adding the last edge to the current acyclic path obtains improvement compared to the tour before reconstruction; (2) All non-adjacent nodes of the current start node have already been selected as target nodes; (3) The number of actions reaches the maximum threshold $M$.
4.3.3 2Opt and Weight Enhancement
After the reconstruction, a tour is constructed by adding the last edge to the reconstructed acyclic path. Subsequently, 2Opt is employed to further optimize the tour, and the search space of 2Opt is also confined to the candidate set of each node, reducing the time complexity from $O(n^{2})$ to $O(nk_{2})$. In a swap operation in 2Opt, two selected edges $x_{i_{1}}$-$x_{j_{1}}$ and $x_{i_{2}}$-$x_{j_{2}}$ are transformed into $x_{i_{1}}$-$x_{j_{2}}$ and $x_{i_{2}}$-$x_{j_{1}}$. For every swap that obtains improvement to the tour, we update $\boldsymbol{Q}$ to enhance the weight of transformed edges as:
$\boldsymbol{Q}_{i,j}\leftarrow\boldsymbol{Q}_{i,j}+\exp{(-\frac{L(\Pi^{new})}{L(\Pi^{pre})})},\ \forall(i,j)\in\{(i_{1},j_{2}),(i_{2},j_{1})\}$ (6)
where $\Pi^{pre}$ is the tour before the swap and $\Pi^{new}$ is the improved tour after the swap. The weight update will guide the next reconstruction, helping the reconstructed path escape local optima while preserving the optimal subpaths. Ultimately, it facilitates the attainment of a better tour in subsequent 2Opt optimization.
5 Experiment
5.1 Datasets
Let TSP-$n$ denote $n$-node TSP instances in uniform distribution. Our mixed-scale training set, abbreviated as TSP-Mix, is derived from a portion of the training data used by Joshi et al. [12] and the detailed features are shown in Section 4.2. As a common setting, the test sets TSP-20/50/100/200/500/1K/10K are taken from part of those used by Fu et al. [6], while TSP-2K/5K are the same as those used by Pan et al. [24]. In detail, the test sets consist of 1024 instances each for TSP-20/50/100; 128 instances each for TSP-200/500/1K; and 16 instances each for TSP-2K/5K/10K. For real-world instances, 78 instances with the number of nodes below 20K are selected from TSPLIB [26].
5.2 Training
In the proposed RsGCN, we set the number of graph convolutional layers $l=6$, the feature dimension $h=128$, and the maximum number of adjacent nodes $k_{1}=\min(50,n)$ for $n$-node instances. The Adam optimizer [13] with a cosine annealing scheduler [21] is employed for training on TSP-Mix. We set $epochs=3$ and $batch\ size=32$, with the learning rate decaying cosinely from 5e-4 to 0. All experiments are conducted on an NVIDIA H20 (96 GB) GPU and an AMD EPYC 9654 (96-Core @ 2.40GHz) CPU. In our configuration, RsGCN only requires approximately 10 minutes per training epoch.
5.3 Baselines
Our proposed solver is compared with: (1) Classical Solvers/Algorithms: Concorde (Applegate et al. 2006), LKH-3 (Helsgaun 2017), 2Opt (Croes 1958); (2) Learning-Based Constructive Solvers: H-TSP (Pan et al. 2023), LEHD (Luo et al. 2023), GLOP (Ye et al. 2024), DRHG (Li et al. 2025); (3) Rule-Based Constructive Solvers: Att-GCN (Fu et al. 2021), SoftDist (Xia et al. 2024), DIFUSCO (Sun et al. 2023), Fast T2T (Li et al. 2024).
All baselines are evaluated on the same test sets and hardware platform as ours. We set $trials=10\text{K}$ for LKH-3 and limit the runtime to a reasonable maximum for Concorde to prevent unacceptable runtime when solving large-scale TSPs. For fairness in dealing with uniform distribution instances, we strive to control the total runtime of LC and RC solvers to be equivalent to or longer than ours. Additional reproduction details on baselines are provided in Appendix C.
5.4 Testing
For Re2Opt, we set the maximum number of threads to 128 to fully utilize the CPU, set the runtime limit to $0.05n$ seconds, and randomly sample the maximum number of reconstruction actions $M\in[10,\min(40,n))$ for $n$-node instances. To account for distributional differences, we set the candidate size $k_{2}$ to 5 and 10 for uniform distribution and real-world instances, respectively.
Three metrics are employed for performance evaluation, including (1) the average TSP tour length; (2) the average optimality gap; (3) the total solution time across all instances for each scale. Herein, the optimality gap of an instance is calculated as $(L/\hat{L})-1$, where $\hat{L}$ denotes the optimal tour length solved by Concorde. We conduct repeated experiments with 5 random seeds, reporting the average metrics for uniform distribution instances and the best metrics for real-world instances in TSPLIB.
5.5 Main Results
Tables 2 and 2, 3 (Table 3 on small-scale TSPs is shown in Appendix D) demonstrate RsGCN’s powerful generalization and state-of-the-art performance. With only 3 epochs of training on TSP-Mix containing instances with up to 100 nodes, RsGCN with Re2Opt as the post-search algorithm can be generalized to 10K-node instances. The proposed Re2Opt can also obtain competitive results without the guidance of RsGCN, validating the efficiency of its search framework. In addition, Figure 6 shows that RsGCN requires the fewest learnable parameters among neural baselines, and RsGCN also requires the fewest training epochs, as shown in Appendix G.
Table 1: Comparison results on TSP-1K/2K/5K/10K. G denotes greedy search. “—” indicates absence of specific reproduction code or available heatmaps. SoftDist generates heatmaps based on distance transform without neural networks. Re2Opt here employs the 5-nearest neighbors as candidates, substituting for 5-hottest neighbors generated by RsGCN to comparatively evaluate RsGCN’s guidance effect on Re2Opt. The two-stage solution time is connected with +. The first-stage times of Att-GCN and DIFUSCO are taken from their publications due to reproducibility issues. Except for Concorde and LKH-3, we mark the smallest and second smallest gaps in red and blue, respectively.
Solver Type TSP-1K TSP-2K TSP-5K TSP-10K
Length Gap(%) Time Length Gap(%) Time Length Gap(%) Time Length Gap(%) Time
Concorde Exact 23.1218 0.0000 1.75m 32.4932 0.0000 1.98m 51.0595 0.0000 5.00m 71.9782 0.0000 10.33m
LKH-3 Heuristic 23.1684 0.2017 17.96s 32.6478 0.4756 29.75s 51.4540 0.7727 2.06m 72.7601 1.0864 7.06m
2Opt Heuristic 24.7392 6.9961 0.12s 34.9866 7.6729 0.49s 55.1501 8.0115 7.87s 77.8939 8.2188 51.88s
H-TSP LC 24.6679 6.6868 36.97s 34.8984 7.4022 9.42s 55.0178 7.7523 20.08s 77.7446 8.0113 38.57s
LEHD LC 23.6907 2.4603 1.48m 34.2119 5.2892 2.15m 59.6522 16.8290 9.59m 90.5505 25.8026 54.45m
GLOP LC 23.8377 3.0962 1.51m 33.6595 3.5893 1.80m 53.1567 4.1073 4.56m 75.0439 4.2592 9.01m
DRHG LC 23.3217 0.8648 1.50m 32.8920 1.2274 2.01m 51.6350 1.1271 5.00m 72.8973 1.2770 9.00m
Att-GCN RC+MCTS 23.6454 2.2650 43.94s + 1.71m 74.3267 3.2627 4.16m + 17.38m
SoftDist MCTS 23.6622 2.3378 0.14s + 1.81m 33.3675 2.6909 0.12s + 3.39m 52.6930 3.1994 0.12s + 8.66m 74.1944 3.0791 0.14s + 17.03m
DIFUSCO RC+MCTS 23.4243 1.3083 0.12s + 1.71m 73.8913 2.6579 0.14s + 17.55m
Fast T2T RC+G+2Opt 23.3122 0.8236 2.95m 32.7940 0.9258 2.03m 51.7409 1.3345 4.35m 72.9450 1.3432 17.00m
Re2Opt Heuristic 23.2756 0.6657 52.38s 32.7452 0.7757 1.68m 51.5199 0.9018 4.18m 72.8742 1.2448 8.37m
[…]
Table 2: Average gaps(%) across different size intervals on TSPLIB instances. The results of other baselines are taken from previous works [19, 17]. Configuration settings are provided below solvers in the header. Complete results on each TSPLIB instance can be found in Appendix E.
Size DIFUSCO T2T Fast T2T BQ LEHD GLOP DRHG Re2Opt RsGCN
Ts=50 Ts=50, Tg=30 Ts=10, Tg=10   bs16 RRC1K more rev. T=1K $k_{2}$=10 $k_{1}$=50, $k_{2}$=10
$<$100 0.73 0.11 0.00 0.49 0.48 0.54 0.48 0.00 0.00
[100, 200) 0.91 0.39 0.25 1.66 0.20 0.79 0.15 0.37 0.00
[200, 500) 2.48 1.42 0.91 1.41 0.38 1.87 0.36 0.90 0.04
[500, 1K) 3.71 1.78 1.19 2.20 1.21 3.28 0.26 0.25 0.13
$\geq$1K 6.68 4.14 7.23 2.09 1.13 0.77
All 2.95 1.59 3.58 0.95 0.72 0.30
5.6 Ablation Study
To validate the effectiveness of the Rescaling Mechanism, we set the following ablation configurations: (1) Rs×2: rescaling adjacent nodes and subgraph edges, i.e., RsGCN; (2) Rs×1: only rescaling adjacent nodes; (3) Rs×0: no rescaling; (4) 5-NN: Re2Opt with the 5-nearest neighbors as candidates. Figure 6 reveals that the Rescaling Mechanism is crucial for improving GCNs’ generalization, where 5-NN acts as the borderline and the gap to 5-NN is presented. Without rescaling edges, Rs×1 shows inferior generalization compared to Rs×2. Even worse, Rs×0 fails to provide effective guidance for TSPs with 1K nodes and more. Due to its quadratic computational complexity, Rs×0 fails to complete inference for TSP-5K/10K within GPU memory. Figure 6 shows that the overall convergence speed follows Rs×2 > Rs×1 > 5-NN > Rs×0, validating the efficient guidance of RsGCN (the y-axis represents the average tour length across 16 instances for each scale, and the x-axis represents the runtime). Furthermore, ablation study on Re2Opt is presented in Appendix F.
Figure 4: Comparison of the number of learnable parameters among neural sovlers.
Figure 5: Comparison of zero-shot generalization.
Figure 6: Comparison of Re2Opt’s Convergence speed on large-scale TSPs.
To visually assess heatmaps’ quality, Figure 8 displays the consistent local regions of the ordered heatmaps generated by Rs×2/×1/×0 on two instances with 500 and 2K nodes, respectively. Some parts of the main diagonal in Rs×0’ heatmap do not obtain ground-truth heats, which easily lead to local optima. While for Rs×1, although the coverage on the main diagonal is enhanced compared to Rs×0, it suffers from the excessive dispersion of heats, which cannot efficiently guide the post-search. Overall, Rs×2’s heatmaps obtain the best coverage and the heat distribution is concentrated around the main diagonal, indicating its effectiveness and superiority.
Furthermore, two metrics are incorporated to evaluate heatmaps: (1) Average Top-k: The average rank of ground-truth nodes’ heats, and (2) Missing Rate in Top-5: The probability that ground-truth nodes are not included in the 5-hottest neighbors. Figure 8 presents the results of these two metrics, where Dist-1 denotes heatmaps obtained by calculating the reciprocal of distance matrices. Figure 8 indicates that Rs×2 achieves superior overall performance, further validating the effectiveness of the Rescaling Mechanism.
Figure 7: Comparison of ordered heatmaps.
Figure 8: Two metrics that reveal the quality of heatmaps (smaller values are preferred).
6 Conclusion
To enhance the generalization capability of GCNs for scalable TSPs, we propose a new RsGCN that incorporates a Rescaling Mechanism.The Rescaling Mechanism includes rescaling adjacent nodes and rescaling subgraph edges to significantly reduce GCN’s sensitivity to scale-dependent features and help RsGCN learn universal patterns across TSPs of various scales. By 3-epoch training on a mixed-scale dataset composed of instances with up to 100 nodes, RsGCN can be successfully generalized to 10K-node instances without any fine-tuning. To efficiently utilize the heatmaps generated by RsGCN, a new post-search algorithm Re2Opt is proposed, which includes a reconstruction process to enhance search diversity and avoid local optima. Extensive experimental results on uniform distribution instances of 9 different scales from 20 to 10K and 78 real-world instances from TSPLIB validate the state-of-the-art performance of the proposed solver. In addition, the training cost and the number of learnable parameters are both significantly lower than those of the existing neural TSP solvers.
Appendix A Ordered Heatmaps
Let $\hat{\boldsymbol{H}}\in\mathbb{R}^{n\times n}$ denote the ordered heatmap for an $n$-node instances. The procedure to obtain an ordered heatmap is described in Algorithm 1. An example of ideal ordered heatmap is shown in Figure 9, in which the neighboring grids along the main diagonal, as well as the grids in the lower-left and upper-right corners of the heatmap, have the largest heat values of 1 while the other grids have the heat value of 0. With such a transformation, we can intuitively investigate the quality of heatmaps by observing the distribution of heat value, i.e., a heatmap similar to the ideal heatmap is preferred.
Input: The original heatmap $\boldsymbol{H}$,
          the optimal tour $\hat{\Pi}$
Output: Ordered heatmap $\hat{\boldsymbol{H}}$
for $i=1$ to $n$ do
       $k\leftarrow\text{index of }i\text{ in }\hat{\Pi}$;
       for $j=1$ to $n$ do
             $h=j+k-1$;
             if $h>n$ then
                   $h=h-n$;
            $\hat{H}_{i,j}=H_{i,\pi_{h}}$;
Algorithm 1 Obtain Ordered Heatmap
Figure 9: The ideal ordered heatmap of TSP-10.
Appendix B Architecture of GCN
Our model adopts Residual Gated Graph Convnets [2, 12]. With regard to the global representation, both $i$ and $j$ index the node set $\mathcal{X}_{n}$ in this section.
Embedding Layer
According to the problem definition, $x_{i}$ denotes the 2-dimensional unscaled coordinates in $\mathcal{X}_{n}$. Let $e_{i,j}$ denote the Euclidean distance between $x_{i}$ and $x_{j}$ ($\in\mathcal{N}_{i}^{k_{1}}$) based on the corresponding rescaled coordinates in $R(\mathcal{N}_{i}^{k_{1}})$. Subsequently, the unscaled coordinate $x_{i}$ and the rescaled edges $e_{i,j}$ are linearly embedded to $h$-dimensional features as follows:
$x_{i}^{0}=\boldsymbol{W}_{1}x_{i}+b_{1}$ (7)
$e_{i,j}^{0}=\boldsymbol{W}_{2}e_{i,j}+b_{2}$ (8)
where $\boldsymbol{W}_{1}\in\mathbb{R}^{2\times h}$ and $\boldsymbol{W}_{2}\in\mathbb{R}^{1\times h}$ are learnable parameter matrices, $b_{1}$ and $b_{2}$ are learnable biases. Such a design of the embedding layer preserves global node distribution while ensuring stable feature representation of subgraphs during GCN’s message aggregation.
Graph Convolutional Layer
Given the embedded features $x_{i}^{0}$ and $e_{i,j}^{0}$, the graph convolutional node features $x_{i}^{\ell}$ and edge features $e_{i,j}^{\ell}$ at layer $\ell\ (\geq 1)$ are defined as follows:
$x_{i}^{\ell}=x_{i}^{\ell-1}+\text{GELU}(\text{LN}(\boldsymbol{W}_{3}^{\ell}x_{i}^{\ell-1}+\sum_{x_{j}\in\mathcal{N}_{i}^{k_{1}}}\sigma(e_{i,j}^{\ell-1})\odot\boldsymbol{W}_{4}^{\ell}x_{j}^{\ell-1}))$ (9)
$e_{i,j}^{\ell}=e_{i,j}^{\ell-1}+\text{GELU}(\text{LN}(\boldsymbol{W}_{5}^{\ell}e_{i,j}^{\ell-1}+\boldsymbol{W}_{6}^{\ell}x_{i}^{\ell-1}+\boldsymbol{W}_{7}^{\ell}x_{j}^{\ell-1}))$ (10)
where $\boldsymbol{W}_{3\sim 7}\in\mathbb{R}^{h\times h}$ are four learnable parameter matrices. LN denotes Layer Normalization, and GELU denotes the Gaussian Error Linear Unit used as the feature activation function. $\sigma$ denotes the Sigmoid function used as the gating activation function. $\odot$ denotes the Hadamard product operation.
Projection Layer
Let $e_{i,j}^{l}$ denote the edge features in the last layer, where $l$ is the maximum number of graph convolutional layers. $e_{i,j}^{l}$ is projected to $H_{i,j}\in[0,1]$ as follows:
$H_{i,j}=\begin{cases}\sigma(\text{MLP}(e_{i,j}^{l})),&\text{if }x_{j}\in\mathcal{N}_{i}^{k_{1}}\text{ and }i\neq j\\ 0,&\text{others}\end{cases}$ (11)
where MLP is a two-layer perceptron with GELU as the activation function. $H_{i,j}$ represents the heat between nodes $i$ and $j$, i.e., the probability that the edge connecting nodes $i$ and $j$ exists in the TSP tour.
Appendix C Reproduction Details on Baselines
For neural solvers, we set the maximum batch size within the GPU memory limits to fully utilize the GPU. We set the maximum number of threads to 128 for CPU-dependent computation solvers Concorde, LKH-3, 2Opt, MCTS, and Re2Opt. For fairness, we adjust the hyperparameters to keep the runtime of baselines approximately the same. However, it remains difficult to ensure complete fairness, as the primary computational resources differ among solvers. LC solvers leverage the GPU’s powerful parallel processing capability to perform larger batch computations (typically covering all instances of each scale in one batch), whereas RC solvers mainly employ the GPU when generating heatmaps, with the post-search algorithms MCTS and Re2Opt relying entirely on the CPU.
Among neural solvers, H-TSP, DRHG, DIFUSCO, and Fast T2T employ fine-tuning strategies. We conduct evaluation using the fine-tuned DRHG for TSP-1K+. H-TSP, DIFUSCO, and Fast T2T are all fine-tuned for multiple specific instance scales. For these three solvers, if no fine-tuned model corresponding to the test scale is available, we use model weights with the fine-tuned scale closest to the test scale for evaluation.
Appendix D Additional Results on TSP-20/50/100/200/500
Table 3: Comparison results on TSP-20/50/100/200/500.
Solver Type TSP-20 TSP-50 TSP-100 TSP-200 TSP-500
Length Gap(%) Time Length Gap(%) Time Length Gap(%) Time Length Gap(%) Time Length Gap(%) Time
Concorde Exact 3.8403 0.0000 1.19s 5.6870 0.0000 3.35s 7.7560 0.0000 18.96s 10.7191 0.0000 7.44s 16.5464 0.0000 31.04s
LKH-3 Heuristic 3.8403 0.0000 28.32s 5.6870 0.0000 56.52s 7.7561 0.0003 55.33s 10.7193 0.0022 7.41s 16.5529 0.0391 10.12s
2Opt Heuristic 3.8459 0.1452 0.04s 5.7812 1.6490 0.07s 8.0124 3.3033 0.15s 11.2332 4.7946 0.08s 17.5688 6.1776 0.03s
H-TSP LC 10.8276 1.0122 6.09s 17.5962 6.3446 17.82s
LEHD LC 3.8403 0.0000 10.00s 5.6870 0.0005 25.00s 7.7572 0.0147 45.00s 10.7309 0.1103 15.00s 16.6603 0.6881 54.00s
GLOP LC 3.8406 0.0078 14.07s 5.6935 0.1143 29.42s 7.7839 0.3597 1.65m 10.7952 0.7099 19.86s 16.9089 2.1907 46.61s
DRHG LC 3.8403 0.0000 10.00s 5.6872 0.0039 25.00s 7.7581 0.0272 45.00s 10.7556 0.3410 15.00s 16.6448 0.5947 50.00s
Att-GCN RC+MCTS 3.8403 0.0000 2.39s + 10.14s 5.6875 0.0090 15.91s + 23.16s 7.7571 0.0137 24.21s + 46.13s 10.7625 0.4035 20.62s + 12.99s 16.7996 1.5308 31.17s + 51.53s
SoftDist MCTS 3.8745 0.8801 0.12s + 9.82s 5.7692 1.4495 0.12s + 23.93s 7.9553 2.5842 0.12s + 45.89s 10.8971 1.6647 0.12s + 13.39s 16.8161 1.6295 0.12s + 51.53s
DIFUSCO RC+MCTS 16.6378 0.5519 3.61m + 51.40s
Fast T2T RC+G+2Opt 3.8416 0.0348 53.00s 5.6876 0.0115 1.00m 7.7587 0.0337 1.23m 10.7445 0.2364 28.00s 16.6461 0.6024 37.00s
Re2Opt Heuristic 3.8403 0.0000 10.31s 5.6874 0.0066 22.40s 7.7626 0.0854 42.14s 10.7473 0.2634 12.46s 16.6309 0.5108 27.15s
[…]
Appendix E Additional Results on TSPLIB
Tables 4 and 5 present the specific optimality gaps on 78 TSPLIB instances, where the results of DIFUSCO, T2T, and Fast T2T are sourced from [19], while those of POMO, BQ, LEHD, and DRHG are extracted from [17]. Due to the distinctive and unusual node distributions of instances fl1577 and fl3795, we activate all neighbors as candidate nodes in Re2Opt’ State Initialization only for solving fl1577 and fl3795. Consequently, the optimality gaps decrease as follows: For Re2Opt, 19.53% $\rightarrow$ 0.71% on fl1577 and 15.23% $\rightarrow$ 1.40% on fl3795; For RsGCN, 8.23% $\rightarrow$ 0.37% on fl1577 and 7.45% $\rightarrow$ 0.97% on fl3795. Note that without such a small adjustment on these two instances, the average optimality gap across all 78 instances obtained by the proposed RsGCN will grow to 0.49, which is still the best compared to other neural solvers.
Table 4: Optimality gaps(%) on 78 TSPLIB instances (OOM refers to GPU out-of-memory).
Instance DIFUSCO T2T Fast T2T POMO BQ LEHD DRHG Re2Opt RsGCN
Ts=50 Ts=50, Tg=30 Ts=10, Tg=10 aug×8 bs16 RRC100 T=1K $k_{2}$=10 $k_{1}$=50, $k_{2}$=10
a280 1.39 1.39 0.10 12.62 0.39 0.30 0.34 0.00 0.00
berlin52 0.00 0.00 0.00 0.04 0.03 0.03 0.03 0.00 0.00
bier127 0.94 0.54 1.50 12.00 0.68 0.01 0.01 0.00 0.00
brd14051 OOM OOM OOM 4.02 2.62 1.91
ch130 0.29 0.06 0.00 0.16 0.13 0.01 0.01 0.00 0.00
ch150 0.57 0.49 0.00 0.53 0.39 0.04 0.04 0.00 0.00
d198 3.32 1.97 0.86 19.89 8.77 0.71 0.26 0.01 0.01
d493 1.81 1.81 1.43 58.91 8.40 0.92 0.31 0.01 0.27
d657 4.86 2.40 0.64 41.14 1.34 0.91 0.21 0.12 0.12
Table 5: Optimality gaps(%) for 78 TSPLIB instances (continued from Table 4).
Instance DIFUSCO T2T Fast T2T POMO BQ LEHD DRHG Re2Opt RsGCN
Ts=50 Ts=50, Tg=30 Ts=10, Tg=10 aug×8 bs16 RRC100 T=1K $k_{2}$=10 $k_{1}$=50, $k_{2}$=10
d1291 77.24 5.97 2.71 2.09 1.74 0.14
d1655 80.99 9.67 5.16 1.57 0.83 1.76
d2103 75.22 15.36 1.22 1.82 0.51 0.17
d15112 OOM OOM OOM 3.41 2.30 2.17
d18512 OOM OOM OOM 3.63 2.28 2.12
eil51 2.82 0.14 0.00 0.83 0.67 0.67 0.67 0.00 0.00
eil76 0.34 0.00 0.00 1.18 1.24 1.18 1.18 0.00 0.00
eil101 0.03 0.00 0.00 1.84 1.78 1.78 1.78 0.00 0.00
fl417 3.30 3.30 2.01 18.51 5.11 2.87 0.49 0.00 0.00
fl1400 47.36 11.60 3.45 1.43 1.67 0.21
fl1577 71.17 14.63 3.71 3.08 0.71 0.37
fl3795 126.86 OOM 7.96 4.61 1.40 0.97
fnl4461 OOM OOM 12.38 1.20 0.89 0.91
gil262 2.18 0.96 0.18 2.99 0.72 0.33 0.33 0.08 0.04
kroA100 0.10 0.00 0.00 1.58 0.02 0.02 0.02 0.00 0.00
kroA150 0.34 0.14 0.00 1.01 0.01 0.00 0.00 0.00 0.00
kroA200 2.28 0.57 0.49 2.93 0.50 0.00 0.00 0.00 0.00
kroB100 2.29 0.74 0.65 0.93 0.01 -0.01 -0.01 0.00 0.00
kroB150 0.30 0.00 0.07 2.10 -0.01 -0.01 -0.01 0.00 0.00
kroB200 2.35 0.92 2.50 2.04 0.22 0.01 0.01 0.00 0.00
kroC100 0.00 0.00 0.00 0.20 0.01 0.01 0.01 0.00 0.00
kroD100 0.07 0.00 0.00 0.80 0.00 0.00 0.00 0.00 0.00
kroE100 3.83 0.27 0.00 1.31 0.07 0.00 0.17 0.00 0.00
lin105 0.00 0.00 0.00 1.31 0.03 0.03 0.03 0.00 0.00
lin318 2.95 1.73 1.21 10.29 0.35 0.03 0.30 0.32 0.00
linhp318 2.17 1.11 0.78 12.11 2.01 1.74 1.69 1.98 1.65
nrw1379 41.52 3.34 8.78 1.41 0.42 0.58
p654 7.49 1.19 1.67 25.58 4.44 2.00 0.03 0.00 0.00
pcb442 2.59 1.70 0.61 18.64 0.95 0.04 0.27 0.35 0.00
pcb1173 45.85 3.95 3.40 0.39 0.83 0.56
pcb3038 63.82 OOM 7.23 1.01 0.69 0.74
pr76 1.12 0.40 0.00 0.14 0.00 0.00 0.00 0.00 0.00
pr107 0.91 0.61 0.62 0.90 13.94 0.00 0.00 0.00 0.00
pr124 1.02 0.60 0.08 0.37 0.08 0.00 0.00 0.00 0.00
pr136 0.19 0.10 0.01 0.87 0.00 0.00 0.00 0.00 0.00
pr144 0.80 0.50 0.39 1.40 0.19 0.09 0.00 7.28 0.00
pr152 1.69 0.83 0.19 0.99 8.21 0.27 0.19 0.18 0.00
pr226 4.22 0.84 0.34 4.46 0.13 0.01 0.01 12.12 0.00
pr264 0.92 0.92 0.73 13.72 0.27 0.01 0.01 0.00 0.00
pr299 1.46 1.46 1.40 14.71 1.62 0.10 0.02 0.00 0.00
pr439 2.73 1.63 0.50 21.55 2.01 0.33 0.12 0.98 0.03
pr1002 43.93 2.94 0.77 0.67 0.37 0.17
pr2392 69.78 7.72 5.31 0.56 0.32 0.56
rat99 0.09 0.09 0.00 1.90 0.68 0.68 0.68 0.00 0.00
rat195 1.48 1.27 0.79 8.15 0.60 0.61 0.57 0.22 0.00
rat575 2.32 1.29 1.43 25.52 0.84 1.01 0.36 0.32 0.21
rat783 3.04 1.88 1.03 33.54 2.91 1.28 0.47 0.42 0.22
rd100 0.08 0.00 0.00 0.01 0.01 0.01 0.01 0.00 0.00
rd400 1.18 0.44 0.08 13.97 0.32 0.02 0.36 0.02 0.00
rl1304 67.70 5.07 1.96 0.79 0.56 0.42
rl1323 68.69 4.41 1.71 1.26 0.20 0.24
rl1889 80.00 7.90 2.90 0.95 1.26 0.14
rl5915 OOM OOM 11.21 1.97 1.31 0.82
rl5934 OOM OOM 11.11 2.69 2.87 1.42
rl11849 OOM OOM 21.43 3.94 3.11 1.79
st70 0.00 0.00 0.00 0.31 0.31 0.31 0.31 0.00 0.00
ts225 4.95 2.24 1.37 4.72 0.00 0.00 0.00 0.00 0.00
tsp225 3.25 1.69 0.81 6.72 -0.43 -1.46 -1.46 -1.40 -1.40
u159 0.82 0.00 0.00 0.95 -0.01 -0.01 -0.01 0.00 0.00
u574 2.50 1.85 0.94 30.83 2.09 0.69 0.24 0.20 0.00
u724 2.05 2.05 1.41 31.66 1.57 0.76 0.27 0.44 0.21
u1060 53.50 7.04 2.80 0.48 0.35 0.39
u1432 38.48 2.70 1.92 0.49 0.69 0.45
u1817 70.51 6.12 4.15 2.05 0.74 0.46
u2152 74.08 5.20 4.90 2.24 0.86 0.56
u2319 26.43 1.33 1.99 0.30 0.50 0.53
usa13509 OOM OOM 34.65 11.82 2.17 1.48
vm1084 48.15 5.93 2.17 0.14 0.25 0.16
vm1748 62.05 6.04 2.61 0.54 0.31 0.17
Average 26.41 2.95 1.59 0.95 0.72 0.30
Appendix F Additional Ablation Study on Re2Opt
Figure 10 shows the average optimal gaps (%) under different combinations of two hyperparameters: the maximum number of actions $M$ and candidate set sizes $k_{2}$. We select 16 instances for each scale. The runtime limit of Re2Opt is set to $0.05n$ seconds for $n$-node instances. The x-axis represents $M$ with different sampling ranges $[10,20]$, $[10,30]$, $[10,40]$, $[10,50]$, and $[10,60]$, while the y-axis represents $k_{2}$ with different values across $[4,7]$. As shown in Figure 10, $k_{2}$ has a more significant impact on Re2Opt. For larger instance scales, a smaller $k_{2}$ tends to help Re2Opt achieve smaller optimality gaps. As a larger $k_{2}$ expands the search space, the efficiency of reconstruction is weakened and the runtime of each 2Opt step is prolonged in dealing with large-scale instances.
Figure 10: Ablation study on the two hyperparameters of Re2Opt.
Appendix G Comparisons of Experimental Configurations
Table 6 presents the configurations of the main training phase for each neural solver, where SL denotes supervised learning and RL denotes reinforcement learning. Our proposed RsGCN requires significantly fewer learnable parameters and training epochs compare to other baselines, substantially reducing the training cost. In GLOP, “×3” indicates that GLOP requires combining 3 neural models with identical parameter sizes during a single solving process, whereas the “+” in H-TSP denotes the combination of two neural models with different parameter scales. Both GLOP and H-TSP employ reinforcement learning, requiring a very large number of training epochs.
Table 6: Comparisons on training configurations for neural solvers.
Solver Learning Paradigm Learnable Parameters Training Epochs
RsGCN SL 0.417M 3
LEHD SL 1.43M 150
DRHG SL 2.65M 100
GLOP RL 1.30M × 3 >500
DIFUSCO SL 5.33M 50
Fast T2T SL 5.33M 50
H-TSP RL 5.34M + 2.51M >500
Att-GCN SL 11.1M 15
Appendix H Visualization of Rescaling Mechanism
Figure 11 visually illustrates the influence of scale-dependent features and the effect of the Rescaling Mechanism. In the normalized square unit space, the distribution of nodes shown in the first row becomes denser as the number of nodes increases. Since nodes are fully connected in TSPs, the number of adjacent nodes for each node is a scale-dependent feature. To rescale adjacent nodes, k-Nearest Neighbor Selection is conducted and each node forms a subgraph containing itself and its top five nearest neighbors in the instance. Note that Figure 11 presents the subgraph of a given node (denoted by orange color) for each instance in the second row. By rescaling adjacent nodes, the subgraphs of nodes from TSP instances with various scales have a uniform number of adjacent nodes, which achieves graph sparsification and helps RsGCN learn universal patterns to enhance generalization capability.
In addition, we observe that the magnitude of edge lengths in the subgraphs still varies in different scales of instances, thereby the length of subgraph edge is another scale-dependent feature. As shown in Figure 11, the average distances between the given node and its neighbors in the three instances are 0.1793, 0.0878, and 0.0206 respectively, which have a significant difference in magnitude and will weaken the generalization capability. Note that the corresponding positions of the subgraphs are denoted by red frames in the distribution of nodes shown in the first row, which intuitively reveal the magnitude difference. Thus, the Rescaling Mechanism further rescales the subgraph edges by Uniform Unit Square Projection to obtain the rescaled subgraph shown in the third row. We can see that the average distances between the given node and its neighbors are in the same magnitude for the three instances, while the topological structure of the subgraphs is preserved. By rescaling subgraph edges, the lengths of subgraph edges across TSPs of various scales are adjusted to the same magnitude, which helps RsGCN generalize the patterns learned from small-scale instances for large-scale instances.
Figure 11: Visualization of the Rescaling Mechanism on TSP-20/200/2000.
Appendix I Limitations and Broader Impacts
Limitations
Our current methodology focuses on solving Symmetric TSP. Future work requires further extensions to broader routing problems, including the Asymmetric TSP (ATSP) and Capacitated Vehicle Routing Problem (CVRP). Our proposed solver still trails behind Concorde since Concorde is an exact solver that can obtain the optimal tour of almost every instance. However, Concorde highly relies on expert-crafted heuristics specifically designed for TSPs. LKH is another solver that also highly relies on expert-crafted heuristics, while the proposed solver can achieve comparable performance to LKH on large-scale TSPs with more than 1K nodes. In addition, the proposed solver significantly outperforms neural baselines.
Broader Impacts
Our work highlights the importance of rescaling scale-dependent features and provides insightful implications for future research to improve the generalization of neural combinatorial optimization (NCO) solvers. Furthermore, our solver’s small parameter scale and low training cost will prompt the community to rethink the practical utility and cost-effectiveness of NCO solvers.